a cart for hauling ore out of a goldmine has a mass of 300 kg. it is movinf at 15 m/s and comes to a stop in 10 s. its change in momentum is (?)
Answer #1

Momentum is given by:

P = m*V

Change in momentum will be given by:

dP = Pf - Pi

dP = m*Vf - m*Vi

m = mass of cart = 300 kg

Vf = final speed = 0 m/sec

Vi = initial speed = 15 m/sec

So,

dP = 300*0 - 300*15

dP = -4500 kg.m/sec

Please Upvote.